In ten years, Zachary will be half his father’s age. If his father was 25 years old when he was born, how old is each now?
Answer by mananth(16946)

let Zachary's age be x
father's age = x+25
...
in ten years their ages will be
Zachary x+10
father x+25+10 =x+35
...
x+10 = 1/2* (x+35)
2(x+10)=x+35
2x+20=x+35
2x-x=35-20
x= 15 son's age
fathers age = 40 years
...
m.ananth@hotmail.ca